The perimeter of a rectangle is 46in. It's length is two more than five times it width. Find its dimensions.
Answer by macston(5194)

.
W=width; L=length=5W+2; P=perimeter=46 in
P=2(L+W) Substitute for L.
46 in=2(5W+2+W) Divide each side by 2.
23 in=6W+2 Subtract 2 from each side
21 in=6W Divide each side by 6
3.5 in=W ANSWER 1: The width is 3.5 inches.
L=5W+2=5(3.5in)+2=17.5in+2=19.5 in ANSWER 2: the length is 19.5 inches.
.
CHECK:
P=2(L+W)
46 in=2(19.5in+3.5in)
46 in=2(23 in)
46 in=46 in